OpenStudy (anonymous):

x+1 over 2= negative 5

jhonyy9 (jhonyy9):

(x+1)/2 = -5 x+1= -10 x= -11

OpenStudy (anonymous):

so x plus one half equal a negative five is negative eleven

jhonyy9 (jhonyy9):

- so if you have wrote there x + (1/2) = -5 than will be 2x+1= -10 2x= -11 x= -11/2

OpenStudy (anonymous):

well it dont have that answer on here

jhonyy9 (jhonyy9):

this is the answer sure

OpenStudy (anonymous):

well it has no solution, negative five and one in a half, negative five and one in a half or four and one half, negative five and one and a half or five and one half

OpenStudy (anonymous):

ok neve mind i got what you was tryin to say